# SCALM: Detecting Bad Practices in Smart Contracts Through LLMs

### v1

@@ 1 Introduction
With the widespread use of blockchain technology, smart contracts have become an important part of the blockchain ecosystem (Sharma, Jindal, and Borah 2023). Smart contracts are computer programs that automatically execute contract terms, controlling assets and operations on the chain. However, due to their public and immutable code, smart contracts have become a significant target for attackers (Li et al. 2023). A total of 464 security incidents occurred in 2023, resulting in losses of up to \$2.486 billion (Zone 2024).
Bad practices refer to poor coding habits or design decisions in the development of smart contracts. Although bad practices may not result in immediate security threats, they could potentially lead to future issues such as performance problems, increased security vulnerabilities, and unpredictable code behavior (Kong, Li, and Li 2024). Additionally, bad practices have hidden economic dangers due to the disruption of regular smart contract activities (Niu et al. 2024).
@@ 3 Method
4:Initialize vector database $VDB$
@@
21:     Retrieve similar vectors $\{\vec{v}_{i}\}$ from $VDB$ using cosine similarity
@@ 3.1 Data Collection and Processing
Our data collection comes from the DAppSCAN database (Zheng et al. 2024b), which includes 39,904 smart contracts with 1,618 SWC weaknesses. First, we extract code snippets with bad practices from these contracts. This extraction process involves static code analysis, using tools to identify and extract code snippets with potential security risks. Then, each code block is converted into a vector through our embedding model and stored in the vector database for subsequent fast matching and retrieval operations.
@@ 3.2 Vector Database
$\text{similarity}(A,B)=\frac{A\cdot B}{||A||_{2}\cdot||B||_{2}}$ (2)
where $A$ and $B$ are two vectors, $A\cdot B$ is their dot product, and $||A|||_{2}$ and $||B|||_{2}$ are their second-paradigms.
@@ 4.1 Experiment Settings
Dataset. In this paper, we use the DAppSCAN dataset (Zheng et al. 2024b) as a knowledge base for detecting bad practices in smart contracts. The dataset contains 39,904 Solidity files with 1,618 SWC weaknesses from 682 real projects. The Smartbugs dataset (Durieux et al. 2020) is also used in the experiment, and a total of 1,894 smart contracts with five types of SWC weaknesses are extracted for comparison experiments. These datasets form the basis of our experimental analysis. Table 1 summarizes the smart contract data used.
@@ 6 Related Work
LLMs have been widely applied and validated for their ability to identify and fix vulnerabilities (Napoli and Gatteschi 2023). In the field of smart contract security, various methods based on LLMs have been proposed, and certain effects have been achieved. Firstly, Boi et al. (Boi, Esposito, and Lee 2024) proposed VulnHunt-GPT, a method that uses the GPT-3 to identify common vulnerabilities in OWASP standards smart contracts.s. Similarly, Xia et al. (Xia et al. 2024) introduced a tool called AuditGPT, which utilizes LLM to automatically and comprehensively verify the ERC rules of smart contracts. They break down large, complex audit processes into small, manageable tasks and design prompts for each ERC rule type to improve audit performance. In terms of fuzz testing, Shou et al. (Shou et al. 2024) presented LLM4FUZZ, which employs LLMs to effectively guide fuzz activities in smart contracts and determine their priorities. This approach can enhance test efficiency and coverage compared to traditional fuzz testing methods.
@@
However, the direct use of pre-trained LLMs is no longer sufficient on some occasions, so many studies have chosen to fine-tune LLMs to meet specific needs. For example, Liu et al. (Liu et al. 2024) proposed PropertyGPT, a system for formal verification of smart contracts by retrieving enhanced property generation. This system utilizes the capability of LLM to automatically generate comprehensive properties of smart contracts, including invariants, preconditions, and postconditions, thus improving the security of smart contracts. On the other hand, Storhaug et al. (Storhaug, Li, and Hu 2023) proposed a novel vulnerability-bound decoding method to reduce the amount of vulnerable code generated by the model. They fine-tuned the LLM to include vulnerability tags when generating code and then prohibited these tags during the decoding process to avoid producing vulnerable codes. Finally, Yang et al. (Yang, Man, and Yue 2024) collected a large number of tagged smart contract vulnerabilities and fine-tuned Llama-2-13B for the automatic detection of vulnerabilities in the decentralized finance (DeFi) domain’s smart contracts (Li et al. 2024b).
@@
In conclusion, while LLMs show great potential for improving smart contract security, the full realization of their potential requires continuous fine-tuning and adaptation of these models in specific contexts.

### v2

@@ 1 Introduction
With the widespread use of blockchain technology, smart contracts have become an important part of the blockchain ecosystem (Sharma et al. 2023). Smart contracts are computer programs that automatically execute contract terms, controlling assets and operations on the chain. However, due to their public and immutable code, smart contracts have become a significant target for attackers (Li et al. 2023). A total of 464 security incidents occurred in 2023, resulting in losses of up to \$2.486 billion (Zone 2024).
Bad practices refer to poor coding habits or design decisions in the development of smart contracts. Although bad practices may not result in immediate security threats, they could potentially lead to future issues such as performance problems, increased security vulnerabilities, and unpredictable code behavior (Kong et al. 2024). Additionally, bad practices have hidden economic dangers due to the disruption of regular smart contract activities (Niu et al. 2024).
@@ 3 Method
4:Initialize vector database VDBVDB
@@
21:  Retrieve similar vectors $\{\vec{v}_{i}\}$ from VDBVDB using cosine similarity
@@ 3.1 Data Collection and Processing
Our data collection comes from the DAppSCAN database (Zheng et al. 2024c), which includes 39,904 smart contracts with 1,618 SWC weaknesses. First, we extract code snippets with bad practices from these contracts. This extraction process involves static code analysis, using tools to identify and extract code snippets with potential security risks. Then, each code block is converted into a vector through our embedding model and stored in the vector database for subsequent fast matching and retrieval operations.
@@ 3.2 Vector Database
[…]
where $A$ and $B$ are two vectors, $A\cdot B$ is their dot product, and $A|||_{2}$ and $B|||_{2}$ are their second-paradigms.
@@ 4.1 Experiment Settings
Dataset. In this paper, we use the DAppSCAN dataset (Zheng et al. 2024c) as a knowledge base for detecting bad practices in smart contracts. The dataset contains 39,904 Solidity files with 1,618 SWC weaknesses from 682 real projects. The Smartbugs dataset (Durieux et al. 2020) is also used in the experiment, and a total of 1,894 smart contracts with five types of SWC weaknesses are extracted for comparison experiments. These datasets form the basis of our experimental analysis. Table 1 summarizes the smart contract data used.
@@ 6 Related Work
Smart contract vulnerability detection has been extensively studied using various approaches. Traditional static analysis tools such as Slither (Feist et al. 2019), Securify (Tsankov et al. 2018), Ethainter (Brent et al. 2020), Vandal (Brent et al. 2018), and eThor (Schneidewind et al. 2020) employ pattern matching and dataflow analysis to identify vulnerabilities. Symbolic execution tools like Mythril (Mueller 2018), Manticore (Mossberg et al. 2019), Osiris (Torres et al. 2018), and Pakala (Palkeo 2025) explore program paths to detect issues such as reentrancy and integer overflow. Fuzzing approaches including Echidna (Grieco et al. 2020) and Harvey (Wüstholz and Christakis 2020) generate test inputs to uncover runtime vulnerabilities.
Recent work has explored deep learning methods for vulnerability detection. Graph neural network (GNN) based approaches (Zhuang et al. 2021; Luo et al. 2024; Zhen et al. 2024; Wang et al. 2024b; Shang et al. 2025) represent smart contracts as graphs to capture structural information. Cross-modality learning (Qian et al. 2023) and contrastive learning (Chen et al. 2024) have also been applied to enhance detection accuracy. Other studies focus on specific vulnerability types, such as reentrancy detection (Song et al. 2025), state-inconsistency bugs (Liu et al. 2025), and NFT-specific defects (Ma et al. 2025). Multimodal learning approaches like SmartInv (Wang et al. 2024a) combine multiple data representations for invariant inference. Deep learning with feature fusion (Li et al. 2025) and bytecode analysis (Xiang et al. 2025; Wang et al. 2023) have also shown promising results.
@@
LLMs have been widely applied and validated for their ability to identify and fix vulnerabilities (Napoli and Gatteschi 2023; Chen et al. 2025). In the field of smart contract security, various methods based on LLMs have been proposed, and certain effects have been achieved. Firstly, Boi et al. (Boi et al. 2024) proposed VulnHunt-GPT, a method that uses the GPT-3 to identify common vulnerabilities in OWASP standards smart contracts.s. Similarly, Xia et al. (Xia et al. 2024) introduced a tool called AuditGPT, which utilizes LLM to automatically and comprehensively verify the ERC rules of smart contracts. They break down large, complex audit processes into small, manageable tasks and design prompts for each ERC rule type to improve audit performance. In terms of fuzz testing, Shou et al. (Shou et al. 2024) presented LLM4FUZZ, which employs LLMs to effectively guide fuzz activities in smart contracts and determine their priorities. This approach can enhance test efficiency and coverage compared to traditional fuzz testing methods.
@@
However, the direct use of pre-trained LLMs is no longer sufficient on some occasions, so many studies have chosen to fine-tune LLMs to meet specific needs. For example, Liu et al. (Liu et al. 2024) proposed PropertyGPT, a system for formal verification of smart contracts by retrieving enhanced property generation. This system utilizes the capability of LLM to automatically generate comprehensive properties of smart contracts, including invariants, preconditions, and postconditions, thus improving the security of smart contracts. On the other hand, Storhaug et al. (Storhaug et al. 2023) proposed a novel vulnerability-bound decoding method to reduce the amount of vulnerable code generated by the model. They fine-tuned the LLM to include vulnerability tags when generating code and then prohibited these tags during the decoding process to avoid producing vulnerable codes. Finally, Yang et al. (Yang et al. 2024) collected a large number of tagged smart contract vulnerabilities and fine-tuned Llama-2-13B for the automatic detection of vulnerabilities in the decentralized finance (DeFi) domain’s smart contracts (Li et al. 2024b).
@@
In conclusion, while LLMs show great potential for improving smart contract security, the full realization of their potential requires continuous fine-tuning and adaptation of these models in specific contexts. Formal verification approaches (Runtime Verification Inc. 2026) and dataset construction efforts (Zheng et al. 2024b) also contribute to advancing the field.